Show that one parsec is equal to 206, 265 AU = 3x10^16 m.
Answer #1

A parsec is distance from observing planet to a star at far distance which creates an arcsecond angle with the solar plane. We assume a distance of earth from sun to be one astronomical unit (AU).

So the distance from earth to the star can be estimated from right angle triangle ABC. AB is 1 AU, angle ACB IS 1 ARCSECOND=1/3600 degree

So distance AC can be= AB/tan(1/3600°)=1AU/tan(1/3600°)=206,264.8 AU or 206,265 AU

1 Au is equal to 1.496×10^8Km

206,265AU= 1.496×206265×10^8=308572.44×10^8=3.0857244×10^13 km

Rounding up the value is 3.08×10^13 km or 3×10^16 m